OpenStudy (anonymous):

A bag contains 3 red marbles, 4 blue marbles, and 5 white marbles. What is the probability of pulling out a white marble, and then a blue marble if none are replaced?

OpenStudy (anonymous):

In your problem given "A bag contains 3 red marbles, 4 blue marbles, and 5 white marbles. What is the probability of pulling out a white marble, and then a blue marble if none are replaced?" If none are replaced the number of marbles available for first picking would be 3R+4B+5W = 3+4+5 = 12 marbles in the first picking one marble is taken out so the availble marbles are 12-1 = 11 now probability of picking white marble: favourable events are 5 as there are 5 marbles of white color and total marblesare 12 so the probability of pulling out a white marble = 5/12 now probability of pulling out a blue marble: favourable events are 4 as there are 4 marbles of blue color total marbles now become 12-1 = 11 so probability of pulling out a blue marble = 4/11 so probability on the whole is (5/12)*(4/11) = 20/132

OpenStudy (anonymous):

So many words, so is 20/132 the answer?